OpenStudy (anonymous):

If e^....... satisfies the equation ........then the value of g(x)=cosx/(cosx+sinx) ,x belongs to 0,pi =?

OpenStudy (anonymous):

\[\LARGE e^{(\sin^2 xsin^4 x+ \sin^6 x..x)\log_e2}\] satisifes equation \[\LARGE x^2-9x+8=0\]

OpenStudy (anonymous):

\[\LARGE g(x)=\frac{\cos x}{\cos x+ \sin x}\]

OpenStudy (anonymous):

\[\LARGE x \epsilon [0,\pi]\]

OpenStudy (anonymous):

@hartnn @ganeshie8 @ikram002p

OpenStudy (anonymous):

a quick thing..can i write lthat e^loge2 as 2? so it becomes 2(sin^2x+sin^4x..)etc?

OpenStudy (ikram002p):

im trying to understant.. so when If e^....... satisfies the equation f(x) that mean if we but e^... in eguation it well be true.. any way we just need to find domain e^... & domain g(x) the interval would be domain e^..cup domain g(x) thats wat i could think of ... maby im wrong

OpenStudy (anonymous):

maybe you are right ..but am i right?

OpenStudy (ikram002p):

mmm not sure the interval u put have "3 pi /4 " which does not belong to the interval of g(x)....cuz primarily=0 at that point.. but how did u figer out this interval ?? share it wid me maby w'll conclude somthing.